# Genetics-based deperturbation analysis for the spin-orbit coupled ${\mathrm{A}}^{1}\Sigma^{+}$ and ${\mathrm{b}}^{3}\Pi_{0^{+}}$ states of LiRb

### v1

@@ I Introduction
Since the turn of the century, the realm of ultracold atoms and molecules has emerged as one of the most fascinating fields for physicists [1] and chemists [2]. Compared to ultracold atoms, ultracold molecules offer a richer internal structure, enhancing application capabilities. The most common techniques used to prepare ultracold molecules in experiments are photoassociation (PA) [3, 4] and magnetoassociation [5]. The latter technique prepares Feshbach molecules [6] that are subsequently transferred to their absolute rovibrational ground state via stimulated Raman adiabatic passage (STIRAP) [7]. It has been productively adopted on bialkali polar molecules including 6Li23Na [8], 6Li40K [9], 23Na39K [10], 23Na40K [11, 12, 13, 14, 15], 23Na87Rb [16, 17], 23Na133Cs [18, 19], 40K87Rb [20], and 87Rb133Cs [21, 22, 23].
As one of the last missing members of the ultracold bialkali polar molecules family, LiRb molecule has several unique features in comparison to the species mentioned above [24, 25]. It exhibits the second largest body-frame dipole moment among fermionic bialkali molecules [26]. Microwave-shielded LiRb molecules promise orders of magnitude higher good-to-bad collision rate ratios compared to existing fermionic polar molecules [27, 28]. As a result, they have the potential to be cooled to unprecedentedly low temperatures for the realization of exotic quantum matters [29]. This is made possible by their large dipole moment and high rotational constant.
@@ III.2 Future prospects
The physical model fitted in the present work are useful for the guidance of preparing ultracold molecules in experiment. As illustrated in Fig. 1, the rovibrational levels in the SO coupled $\rm A^{1}\Sigma^{+}$ and $\rm b^{3}\Pi_{0^{+}}$ states could be chosen as the intermediate state for STIRAP. With the fitted PECs and SO coupling, one could calculate the TDM matrix elements, which are helpful in predicting feasible intermediate state for STIRAP transfer.

### v2

@@ I Introduction
Since the turn of the century, the realm of ultracold atoms and molecules has emerged as one of the most fascinating fields for physicists [1] and chemists [2]. Compared to ultracold atoms, ultracold molecules offer a richer internal structure, enhancing application capabilities. The most common techniques used to prepare ultracold molecules in experiments are photoassociation (PA) [3, 4] and magnetoassociation (MA) [5]. The latter technique prepares Feshbach molecules [6] that are subsequently transferred to their absolute rovibrational ground state via stimulated Raman adiabatic passage (STIRAP) [7]. It has been productively adopted on bialkali polar molecules including 6Li23Na [8], 6Li40K [9], 23Na39K [10], 23Na40K [11, 12, 13, 14, 15], 23Na87Rb [16, 17], 23Na133Cs [18, 19], 40K87Rb [20], and 87Rb133Cs [21, 22, 23].
However, ultracold LiRb molecules, especially the fermionic species, have not been deterministically prepared in the absolute rovibrational ground state via MA and subsequently STIRAP. LiRb molecule has several unique features in comparison to the species mentioned above [24, 25]. It exhibits the second largest body-frame dipole moment among fermionic bialkali molecules [26]. Microwave-shielded LiRb molecules promise orders of magnitude higher good-to-bad collision rate ratios compared to existing fermionic polar molecules [27, 28]. As a result, they have the potential to be cooled to unprecedentedly low temperatures for the realization of exotic quantum matters [29]. This is made possible by their large dipole moment and high rotational constant.
@@ III.2 Future prospects
The physical model fitted in this work are useful for finding pathway to produce ultracold molecules in the absolute rovibrational ground state. In previous work, Stevenson $et\ al.$ proposed several STIRAP pathways to transfer bosonic 7Li85Rb molecules to the absolute rovibrational ground state. In [40, 42], they produced 7Li85Rb molecules in $\rm X^{1}\Sigma^{+}$ $v=43$ level through PA followed by spontaneous decay, and suggested using $\rm C^{1}\Sigma^{+}$ $v=22$ level as intermediate state for STIRAP to transfer them to $\rm X^{1}\Sigma^{+}$ $v=0$ level. In [41], they also proposed a STIRAP pathway to transfer Feshbach 7Li85Rb molecules to the absolute rovibrational ground state, with the SO coupled $\rm D^{1}\Pi$ and $\rm d^{3}\Pi$ states as intermediate state. Besides, as illustrated in Fig. 1, the rovibrational levels in the SO coupled $\rm A^{1}\Sigma^{+}$ and $\rm b^{3}\Pi_{0^{+}}$ states could also be chosen as the intermediate state for STIRAP. One possible STIRAP pathway has been recommended in [24], which chooses $\rm A^{1}\Sigma^{+}$ $v=5$ level or $\rm b^{3}\Pi_{0^{+}}$ $v=12$ level as intermediate state. However, the STIRAP pathway proposed in [24] is not based on deperturbation analysis, and could not be accurately generalised to other isotopes, such as 6Li87Rb molecule. With the PECs and SO coupling fitted in this work, we could calculate the TDM matrix elements, which are particularly helpful in predicting feasible intermediate state for STIRAP transfer.